Date: Wed, 25 Apr 2001 05:15:47 -0600 From: "s. a. snedden" Subject: 2.5-m ObsLog 20010424 (MJD 52024) APACHE POINT OBSERVATORY SDSS 2.5M OBSERVING LOG Tuesday, April 24, 2001 (MJD 52024) OBSERVING TEAM: =============== Steph Snedden (APO) Dan Long (APO) Anders Jorgenson (LANL) - PT VISITORS: ========= THE PLAN: ========= Science. Imaging if conditions warrant, otherwise spectroscopy. SYNOPSIS FOR NIGHT: =================== Spectroscopy. Three plates declared done. GENERAL COMMENTS (see Problems below for details): ================================================== Although it was mostly clear before sunset, there were clouds approaching from the south and a forecast for an east wind. We estimated that the clouds would reach us at about 22:00, so we set up for spectroscopy. Sure enough, at sunset the wind shifted East and gave us poor seeing. During afternoon checkout, Steph tried to duplicate the problem she saw yesterday during checkout. After slewing to 6 degrees, she then used the MCP "j" to lower the telescope to 5.5 deg. Yesterday while doing this, the speed went very high and the MCP froze with the telescope at 3.5 degrees. The MCP had to be rebooted. Today, the problem could not be duplicated. We speculate that this could have been caused by a conflict between a telrun session and the MCP, or perhaps a spurious high value got passed to the MCP, but we really don't know. French warned us about fibers dropping out easily on cartridge 2. Although Dan thought we could fix this with 640 tiny pieces of duct tape or velcro ;>), we decided to restrict ourselves to just inspecting the fibers while the cartridge was in the bay. We noted no dropped fibers at that time. Unfortunately, we didn't get to observe this plate tonight, so we don't yet know if a ride over the grate will drop any fibers. When we do observe it, we want to check the fibers before it is mounted. Seeing was predictably bad, so our first plate of the night, 516 (cartridge 4), took a total exposure of 5400 s to declare done. Although the seeing on the next plate (522 on cartridge 3) appeared to be somewhat worse by the guider, the efficiency was quite a bit higher (about 60%). This makes us wonder if the guide fiber ends at the guider end have gotten misaligned, and if so, can they be realigned easily? I understand from Dan they terminate in a block. Can this be adjusted? Plate 522 was declared done with a total exposure time of 3000s. The next plate was 536 (cartridge 7). It was declared done with a total exposure of 3600 s. Next up was Cartridge 1, plate 592. We exposed for 5400 s, but were unable to declare this plate done because of increasing cloud cover. DATA ACQUISITION (see IOP/SOP output below): ============================================ IMAGING: -------- Run Stripe Strip Flavor Lambda Last Frame # Begin End ------------------------------------------------------------------ 2268 100 O ignore -83.07 -80.79 29 doghouse bias SPECTROSCOPY: ------------- OFFSETS: -------- getclock Calib offsets Scale Time Ct Plate time Az El Rot Factor 20:35 4 517 988166102 -0.000136 0.005558 0.007483 1.000280 23:13 3 522 988175622 -0.002348 -0.001822 0.055040 0.999990 00:46 7 536 988181180 -0.002502 0.002169 -0.010150 1.000050 02:50 1 592 988188652 -0.005330 -0.000829 0.001907 0.999760 FOCUS LOG: ---------- setmir piston Temp Wind Time Inst scale M1 M2 Foc Az Alt C MPH Dir filt fwhm " 17:20 East door open, West door cracked, Exhaust and cone fans on 19:25 Enclosure opened 20:35 4 517 1.00028 -2765 -1039 50 55 44.7 5.4 18 79 - 2.4 21:29 4 517 1.00004 -395 895 50 39 53.1 4.2 18 79 - 3.2 22:06 4 517 0.99999 99 1292 50 25 57.2 4.0 15 78 - 2.7 22:35 4 517 0.99996 395 1532 50 11 59.1 3.8 14 78 - 2.5 23:17 3 522 0.99999 99 1236 0 10 59.1 3.8 14 82 - 2.9 00:12 3 522 0.99999 99 1238 0 -17 58.4 3.9 11 81 - 2.8 00:46 7 536 1.00005 -494 675 -75 20 57.5 4.2 10 58 - 2.2 01:22 7 536 1.00005 -494 671 -75 3 59.0 5.4 5 93 - 2.4 02:51 1 592 0.99976 2369 2965 -150 -18 59.4 6.0 7 121 - 1.7 03:33 1 592 0.99979 2073 2731 -150 -36 55.4 5.9 9 117 - 1.9 04:27 1 592 0.99979 2073 2756 -150 -54 47.3 6.2 10 128 - 2.2 TAPES: ------ Ship to FNAL: JL3949 Stays at APO: JL3950 MIRROR NUMBERS: =============== PRIMARY: -------- Scale: 1.000000 MIGS TONIGHT NOMINAL Axial A: 0.0280 0.0260 Axial B: -0.0320 -0.0350 Axial C: 0.0420 0.0350 Trans D: 3.7580 3.7550 Lateral E: 1.9070 1.9065 Lateral F: 0.4085 0.4080 GALILS Commanded: 4890. -2755. 465. -6642. 8737. 8322. Actual: 4900. -2750. 450. -6650. 8750. 8300. SETMIR VALUES primDesOrient: 0.00 -4.40 10.70 586.50 240.70 primOrient: 0.00 -4.56 10.53 586.58 240.03 SECONDARY: ---------- Focus: 0.0 Air temp: 11.4 C Altitude: 30.02 MIGS TONIGHT NOMINAL Axial A: -1.2920 -1.2930 Axial B: -1.3050 -1.3070 Axial C: -1.3030 -1.3050 Trans D: 0.1810 0.1900 GALILS Commanded: 1583837. 1583837. 1583837. 2595. 3795. Actual: 1583970. 1583832. 1583800. 2600. 3800. SETMIR VALUES secDesOrient: 1257.00 0.00 0.00 0.00 -175.00 secOrient: 1257.00 0.01 -0.02 -0.01 -175.13 WEATHER SYNOPSIS: ================= Mostly clear until about 01:30, then partly-to-mostly cloudy. Poor-to-moderate seeing. WEATHER DATA: ============= Wind Time Temp F Dewp F MPH Direction Enchantment DIMM Sky Precip. 16:01 57 15 14 248 (WSW) 573 - mclear none 17:01 59 16 15 241 (WSW) 646 - mclear none 18:01 56 17 16 204 (SSW) 597 - mclear none 19:01 49 16 8 111 (ESE) 576 - mclear none 20:01 44 17 11 82 (E) 587 - mclear none 21:01 41 16 16 81 (E) 593 - mclear none 22:01 39 16 15 73 (ENE) 636 - mclear none 23:01 39 14 13 75 (ENE) 728 - mclear none 00:01 39 14 12 79 (ENE) 639 - mclear none 01:01 41 14 9 81 (E) 704 - pcl none 02:01 41 14 2 113 (ESE) 748 - pcl none 03:01 43 14 10 128 (SE) 675 - pcl none 04:01 42 14 8 110 (ESE) 754 - pcl none 05:01 43 12 9 110 (ESE) 766 - mcl none [pmt]cl = partly, mostly, totally cloudy LN2 SCHEDULE: ============= Imager on autofill Spectrographs on autofill Spectro dewar weight at 05:00: 105 lb. Imager dewar swapped at 05:00 Imager dewar weight at 05:00: 219 lb. TELESCOPE STATUS: ================= Opened doors: 17:00 Blower on: 17:20 Cone fans on: 17:30 Opened enclosure: 19:35 Closed enclosure: 04:50 Blower & cone fans off: 05:00 Stowed at 30 deg alt with cartridge 1 mounted. Counterweights at 270. JUREK'S PLANTS STATUS: ====================== Still have plenty of water Plant 1: 7.0 cm Plant 2: 5.0 cm Plant 3: 1.3 cm INTERLOCKS: =========== None bypassed SOFTWARE USED: ============== IOP/SOP: v3_41_5a Watcher: v2_8_16 MCP: v4_6_5 AstroDa: 14_36 TCC: 2.5.1a TPM: v2_3_3 tccmon: v1_6 sdssProcedures: v1_15 PROBLEMS IN DETAIL: =================== DATA LOGS FROM IOP/SOP: =======================